OpenStudy (anonymous):

The formula is given by (2x^3 +3x^2 +10x-9/(x-1) when x<-1 and by the formula 2x^2-1x+a when -1> or equal to -1 what does a need to be to make the function continuous at -1

OpenStudy (accessdenied):

What is your definition of something being continuous? Just, how you think about it and not necessarily the textbook definition.

OpenStudy (jdoe0001):

have you covered piecewise functions yet?

OpenStudy (anonymous):

that's what im confused about, I know what a continuous function is. This function is saying that it is discontinuous at -1 , but it think Im confused about what they are asking

OpenStudy (jdoe0001):

one of the functions has a variable "a" and depending on that variable's value, THAT FUNCTION will meet THE OTHER FUNCTION

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

still confused?

OpenStudy (anonymous):

no I perfectly understand that, im just a little confused on how to go about solving. I dont thing im connecting the concept with the applicability of the problem

OpenStudy (jdoe0001):

\(\bf f(x)= \begin{cases} \cfrac{2x^3 +3x^2 +10x-9}{x-1} & x< -1 \\ \quad \\ \bf 2x^2-1x+{\color{red}{ a}} & x \ge -1 \end{cases} \\ \quad \\ when\quad x={\color{red}{ a}}\quad \cfrac{2x^3 +3x^2 +10x-9}{x-1}=\textit{some value} \\ \quad \\ when\quad x={\color{red}{ a}}\quad 2x^2-1x+{\color{red}{ a}}=\textit{same exact value as above} \\ \quad \\ \textit{thus we can say} \\ \quad \\ \cfrac{2x^3 +3x^2 +10x-9}{x-1}=2x^2-1x+{\color{red}{ a}}\)

OpenStudy (anonymous):

thank u!

OpenStudy (jdoe0001):

yw